# Split Gibbs Discrete Diffusion Posterior Sampling

### v1

@@ 4.6 Ablation Studies
Effectiveness of the annealing noise schedule. We now investigate how the choice of noise schedule $\{\rho_{k}\}_{k=0}^{K}$ affects the performance of SG-DPS. In our main experiments, we use a geometric noise schedule from (Lou et al., 2024), defined as $\rho_{k}=\rho_{\min}^{k/K}\rho_{\max}^{1-k/K}$. We argue that an annealing noise schedule $\{\rho_{k}\}_{k=0}^{K}$ that gradually decays to $\rho_{\min}\approx 0$ allows SG-DPS to converge faster and more accurately to the posterior distribution.
@@
As shown in Figure 5, running SG-DPS with the annealing noise schedule achieves the highest PSNR and also the fastest convergence speed. Moreover, when $\rho$ is large, SG-DPS converges quickly to a poor solution, whereas when $\rho$ is small, it converges slowly to an accurate solution. This finding aligns with our intuition that SG-DPS is most accurate with $\rho\to 0$, while sampling with $\rho>0$ regularizes and simplifies the sampling problem.
@@
This trade-off between convergence speed and accuracy results in a non-monotonic relationship between PSNR and fixed noise levels: PSNR first increases and then decreases with $\rho$. Therefore, using an annealing noise schedule helps provide both fast initial convergence and accurate posterior sampling as $\rho_{k}$ approaches 0.
Figure 5: Convergence speed of different noise schedules. We compare our noise annealing scheduler $\{\rho_{k}\}_{k=1}^{K}$ (orange) to fixed noise schedules of various noise levels on the MNIST AND task. $y$-axis is the PSNR of variables ${\mathbf{x}}^{(k)}$ with respect to the ground truth. Our noise schedule converges faster than schedulers with fixed noises.
@@
Time efficiency. We run SG-DPS with different configurations to evaluate its performance under different computing budgets. We control the number of iterations, $K$, and the number of unconditional sampling steps $T$ in each prior sampling step, so the number of function evaluations (NFE) of the discrete diffusion model is $KT$. We evaluate SG-DPS with NFE ranging from $100$ to $2$k. The detailed configuration is specified in Section B.1.
@@
We present a plot in Figure 6 showing the tradeoff between sample quality and computing budgets in terms of NFEs.  SG-DPS is able to generate high-quality samples with a reasonable budget comparable to an unconditional generation.
@@
Comparison to sampling without prior. We evaluate the value of discrete diffusion priors in solving discrete inverse problems by comparing our method to the Metropolis-Hasting algorithm without diffusion prior. We pick the AND problem on discretized MNIST as an example, in which we choose $\gamma D$ random pairs of pixels and compute their AND values as measurement ${\mathbf{y}}$. Figure 7 shows the reconstructed samples of SG-DPS and direct MCMC sampling under different levels of measurement sparsity. As demonstrated in the figure, the discrete diffusion model complements the missing information when the measurement is sparse (e.g., $\gamma=2,4,8$), underscoring the importance of discrete diffusion priors in SG-DPS for solving discrete inverse problems.
@@
Figure 7: Comparison of SG-DPS and direct sampling without data prior. The PSNR values are computed with 10 test samples. SG-DPS replenishes the missing information when the measurement is sparse.
5 Conclusion
We introduce SG-DPS, a principled discrete diffusion posterior sampling algorithm based on the split Gibbs sampler, extending plug-and-play diffusion methods from Euclidean spaces to discrete-state spaces. Our algorithm iteratively alternates between prior and likelihood sampling steps with increasing regularization. By carefully designing the regularization potential function, we establish a connection between prior sampling steps and discrete diffusion samplers, integrating discrete diffusion models into the split Gibbs sampling framework. We evaluate our method on diverse tasks from various domains, including solving inverse problems and sequence generation guided by reward functions, where it significantly outperforms existing baselines.
Acknowledgments
We would like to thank the Kortschak scholarship for supporting Wenda Chu at Caltech.
Appendix A Experimental details
A.1 Pretrained Models
@@
We learn prior distributions for each dataset using SEDD (Lou et al., 2024) discrete diffusion models. We use the SEDD small architecture with around 90M parameters for all experiments, and the models are trained with AdamW (Loshchilov & Hutter, 2019) with batch size $32$ and a learning rate of $3\times 10^{-4}$.
A.2 Baseline Methods
A.2.1 DPS
@@
DPS (Chung et al., 2023) is designed to solve general inverse problems with a pretrained (continuous) diffusion model. It performs posterior sampling from $p({\mathbf{x}}|{\mathbf{y}})$ by modifying the reverse SDE
$\mathrm{d}x_{t}$ $=-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\mathbf{y}})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}t$ (24)
@@
However, DPS is not directly applicable to inverse problems in discrete-state spaces since propagating gradients through ${\bm{G}}$ and ${\bm{D}}_{\theta}$ in Equation 26 is impossible. Therefore, we consider the counterpart of DPS in discrete spaces. We modify the continuous-time Markov chain of the discrete diffusion model by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}=\tilde{\bm{Q}}_{T-t}p_{T-t},$ (27)
@@
$\tilde{\bm{Q}}_{t},\tilde{\mathbf{x}}{\mathbf{x}}=\frac{p_{t}(\tilde{\mathbf{x}}|{\mathbf{y}})}{p_{t}({\mathbf{x}}|{\mathbf{y}})}=\frac{p_{t}(\tilde{\mathbf{x}})}{p_{t}({\mathbf{x}})}\frac{p_{t}({\mathbf{y}}|\tilde{\mathbf{x}})}{p_{t}({\mathbf{y}}|{\mathbf{x}})}={\bm{Q}}_{t},\tilde{\mathbf{x}}{\mathbf{x}}\cdot\frac{p_{t}({\mathbf{y}}|\tilde{\mathbf{x}})}{p_{t}({\mathbf{y}}|{\mathbf{x}})}$ (28)
Similar ideas are applied to classifier guidance for discrete diffusion models (Nisonoff et al., 2024), where the matrix ${\bm{R}}_{t}^{\mathbf{y}}=[\frac{p_{t}({\mathbf{y}}|\tilde{\mathbf{x}})}{p_{t}({\mathbf{y}}|{\mathbf{x}})}]_{{\mathbf{x}}\tilde{\mathbf{x}}}$ is called a guidance rate matrix. We compute ${\bm{R}}_{t}^{\mathbf{y}}$ at ${\mathbf{x}}$-column by enumerating every neighboring $\tilde{\mathbf{x}}$ and calculating $\frac{p_{t}({\mathbf{y}}|\tilde{\mathbf{x}})}{p_{t}({\mathbf{y}}|{\mathbf{x}})}$ for each $\tilde{\mathbf{x}}$. The discrete version of DPS can be summarized by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}=\bar{\bm{Q}}_{T-t}{\bm{R}}_{t}^{\mathbf{y}}p_{T-t}.$ (29)
However, the discrete version of DPS is very time-consuming, especially when the vocabulary size is large, for it enumerates $(N-1)\times D$ number of neighboring $\tilde{x}$ when computing ${\bm{R}}_{t}^{\mathbf{y}}$. We find it slow for binary MNIST ($N=2$) and DNA generation tasks $(N=5)$ but unaffordable for monophonic music generation $(N=129)$.
A.2.2 SVDD
SVDD (Li et al., 2024) aims to sample from the distribution $p^{(\alpha)}({\mathbf{x}}_{0})\ \ \propto\ \ p({\mathbf{x}}_{0})\exp(r({\mathbf{x}}_{0})/\alpha)$, which is equivalent to the regularized MDP problem:
$p^{(\alpha)}({\mathbf{x}}_{0})=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{{\mathbf{x}}_{0}\sim\pi}r({\mathbf{x}}_{0})-\alpha\mathrm{KL}(p_{0}\|\pi)$ (30)
@@ A.2.2 SVDD
$v_{t}({\mathbf{x}}_{t})=\alpha\log\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})}[\exp(r({\mathbf{x}}_{0})/\alpha)]$ (31)
@@
$p_{t}^{\star}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})\ \ \propto\ \ p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})\exp(v_{t}({\mathbf{x}}_{t})/\alpha).$ (32)
In time step $t$, SVDD samples a batch of $M$ particles from the unconditional distribution $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})$, and conduct importance sampling according to $\exp(v_{t}({\mathbf{x}}_{t})/\alpha)$.
Although (Li et al., 2024) is initially designed for guided diffusion generation; it also applies to solving inverse problems by carefully choosing reward functions. We consider $r({\mathbf{x}})=-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|_{0}/\sigma_{\mathbf{y}}$ and $\alpha=1$, so that it samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. As recommended in (Li et al., 2024), we choose $\alpha=0$ in practice, so the importance sampling reduces to finding the particle with the maximal value in each iteration. We use SVDD-PM, a training-free method provided by (Li et al., 2024) in our experiments. It approximates the value function by $v_{t}({\mathbf{x}}_{t})=r(\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t}))$, where $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ is an approximation of $\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]$. In practice, we find that approximating $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ by a few-step discrete sampler achieves slightly better results.
A.3 Hyperparameters
We use an annealing noise schedule of $\rho_{k}=\rho_{\min}^{k/K}\rho_{\max}^{1-k/K}$ with $\rho_{\min}=10^{-4}$ and $\rho_{\max}=20$. We run SG-DPS for $K$ iterations. In each likelihood sampling step, we run Metropolis-Hasting for $T$ steps, while in each prior sampling step, we run a few-step Euler discrete diffusion sampler with $L$ steps. The hyperparameters used for each experiment are listed in Table 5. We also include the dimension of data spaces $\mathcal{X}^{D}$ for each experiment in Table 5, where $|\mathcal{X}|=N$.
Synthetic MNIST XOR MNIST AND Music infilling Guided DNA generation
$T$ 10 2000 5000 5000 200
$K$ 10 50 100 100 50
$L$ 20 20 20 20 20
$D$ 2/5/10 1024 1024 256 200
$N$ 50 2 2 129 5

### v2

@@ Appendix B Ablation studies
We investigate how the choice of noise schedule $\{\eta_{k}\}_{k=0}^{K}$ affects the performance of SGDD. In our main experiments, we use a geometric noise schedule from [29], defined as $\eta_{k}=\eta_{\min}^{k/K}\eta_{\max}^{1-k/K}$. We argue that an annealing noise schedule $\{\eta_{k}\}_{k=0}^{K}$ that gradually decays to $\eta_{\min}\approx 0$ allows SGDD to converge faster and more accurately to the posterior distribution.
@@
As shown in Fig. 5, running SGDD with the annealing noise schedule achieves the highest PSNR and also the fastest convergence speed. Moreover, when $\eta$ is large, SGDD converges quickly to a poor solution, whereas when $\eta$ is small, it converges slowly to an accurate solution. This finding aligns with our intuition that SGDD is most accurate with $\eta\to 0$, while sampling with $\eta>0$ regularizes and simplifies the sampling problem.
@@
This trade-off between convergence speed and accuracy results in a non-monotonic relationship between PSNR and fixed noise levels: PSNR first increases and then decreases with $\eta$. Therefore, using an annealing noise schedule helps provide both fast initial convergence and accurate posterior sampling as $\eta_{k}$ approaches 0.
Theoretical insights. Since Eq. 41 holds when $\eta\to 0$, a natural question is why we cannot fix a small $\eta$ to begin with.
The answer to this question is contained in Theorem 1. When fixing a small $\eta$, we also fix a small $t^{*}$ as $\sigma(t^{*})=\eta$. Then the transition matrix ${\bm{Q}}_{\tau}$ ($\tau\in[0,t^{*}]$) will be ill-posed, in the sense that ${\bm{Q}}_{\tau}^{[i,j]}$ is either very large or very close to 0. In this case, the relative Fisher information $\mathsf{FI}_{{\bm{Q}}_{\tau}}(\mu_{\tau}\|\pi_{\tau})$ becomes a poor criterion of measuring the divergence between $\mu_{\tau}$ and $\pi_{\tau}$. To see this, we note that the spectral gap $\gamma({\bm{Q}}_{\tau})$ has
$\gamma({\bm{Q}}_{\tau})=\min_{\lambda\neq 0}-\mathrm{Re}\lambda({\bm{Q}}_{\tau})\geq N\epsilon,$ (42)
where $N=|\mathcal{X}|$, if ${\bm{Q}}_{\tau}^{i,j}>\epsilon\geq 0$ for all $i\neq j$. By the logarithmic Sobolev inequality [3], we have that
$\mathsf{FI}_{{\bm{Q}}_{\tau}}(\mu_{\tau}\|\pi_{\tau})\geq\frac{n\min_{i\neq j}{\bm{Q}}_{\tau}^{[i,j]}}{2}\cdot\mathsf{KL}(\mu_{\tau}\|\pi_{\tau})$ (43)
Ignoring the errors in the score function and discretization, the KL divergence converges linearly, i.e.,
$\mathsf{KL}(\mu_{t}\|\pi_{t})\leq\exp\left(-\frac{n}{4}\int_{0}^{t}\min_{i\neq j}{\bm{Q}}_{\tau}^{[i,j]}\mathrm{d}\tau\right)\mathsf{KL}(\mu_{0}\|\pi_{0}).$ (44)
When $\eta_{k}$ is large, $\int_{0}^{t^{*}}\min_{i\neq j}{\bm{Q}}_{\tau}^{[i,j]}\mathrm{d}\tau$ is considerably large, because the probability ratio $p_{t}({\mathbf{x}})$ is being smoothed by the uniform transition kernel, which makes ${\bm{Q}}_{t}^{[i,j]}$ large. However, as discussed above, $\int_{0}^{t^{*}}{\bm{Q}}_{\tau}^{[i,j]}\mathrm{d}\tau$ can be very close to 0 when $\eta_{k}$ is small, which makes the convergence slow.
B.2 Time efficiency.
@@
We run SGDD with different configurations to evaluate its performance under different computing budgets. We control the number of iterations, $K$, and the number of unconditional sampling steps $T$ in each prior sampling step, so the number of function evaluations (NFE) of the discrete diffusion model is $KT$. We evaluate SGDD with NFEs ranging from $100$ to $2$k. The detailed configuration is specified in Table 6.
@@
We present a plot in Fig. 5 showing the tradeoff between sample quality and computing budgets in terms of NFEs. SGDD is able to generate high-quality samples with a reasonable budget comparable to unconditional generation. Note that we calculate the NFEs for SVDD and SMC with the number of Monte Carlo samples $\mathrm{mc}=1$ and the number of Euler steps $H=1$, while some results in our experiments are done with larger $\mathrm{mc}$ and $H$ for better performance. We report both the number of total NFEs and the number of sequential NFEs that decide the cost of the algorithm when it is fully parallelized.
Table 6: Configurations of SGDD and baseline methods. For DPS and SVDD-PM, we calculate both the total NFE and the sequential NFE. The numbers are collected from the MNIST XOR task as an example. The runtime is amortized on a batch size of 10 samples with one NVIDIA A100 GPU.
Configuration ##\## Iterations Prior sampling steps NFE (total/sequential) Runtime (sec/sample)
DPS $100$ $102400/100$ 820
SVDD-PM ($M=20$) $128$ $2560/128$ 16
SMC $(M=20)$ $128$ $5120/256$ 21
SGDD-$50$ $25$ $2$ $50$ 4
SGDD-$100$ $25$ $4$ $100$ 6
SGDD-$200$ $40$ $5$ $200$ 8
SGDD-$400$ $40$ $10$ $400$ 10
SGDD-$1$k $50$ $20$ $1000$ 13
SGDD-$2$k $100$ $20$ $0020$
B.3 Comparison to sampling without prior.
@@
We evaluate the value of discrete diffusion priors in solving discrete inverse problems by comparing our method to the Metropolis-Hastings algorithm without a diffusion prior. We pick the AND problem on discretized MNIST as an example, in which we choose $\gamma D$ random pairs of pixels and compute their AND values as measurement ${\mathbf{y}}$. Fig. 6 shows the reconstructed samples of SGDD and direct MCMC sampling under different levels of measurement sparsity. As demonstrated in the figure, the discrete diffusion model complements the missing information when the measurement is sparse (e.g., $\gamma=2,4,8$), underscoring the importance of discrete diffusion priors in SGDD for solving discrete inverse problems.
@@
Figure 6: Comparison of SGDD and direct sampling without data prior. The PSNR values are computed with 10 test samples. SGDD replenishes the missing information when the measurement is sparse.
Appendix C Experimental details
C.1 Pretrained Models
@@
We learn prior distributions for each dataset using SEDD [29] discrete diffusion models. We use the SEDD small architecture with around 90M parameters for all experiments, and the models are trained with AdamW [28] with batch size $32$ and a learning rate of $3\times 10^{-4}$.
C.2 Baseline Methods
C.2.1 DPS
@@
DPS [7] is designed to solve general inverse problems with a pretrained (continuous) diffusion model. It performs posterior sampling from $p({\mathbf{x}}|{\mathbf{y}})$ by modifying the reverse SDE
$\mathrm{d}{\mathbf{x}}_{t}$ $=-2\dot{\sigma}_{t}\sigma_{t}\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\mathbf{y}})\mathrm{d}t+\sqrt{2\dot{\sigma}_{t}\sigma_{t}}\mathrm{d}t$ (45)
@@
However, DPS is not directly applicable to inverse problems in discrete-state spaces since propagating gradients through ${\bm{G}}$ and ${\bm{D}}_{\theta}$ in Eq. 47 is impossible. Therefore, we consider the counterpart of DPS in discrete spaces. We modify the continuous-time Markov chain of the discrete diffusion model by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}={\bm{Q}}_{t}^{\mathbf{y}}p_{T-t},$ (48)
@@
${{\bm{Q}}}_{t}^{{\mathbf{y}}\ [i,j]}=\frac{p_{T-t}({\mathbf{x}}_{i}|{\mathbf{y}})}{p_{T-t}({\mathbf{x}}_{j}|{\mathbf{y}})}=\frac{p_{T-t}({\mathbf{x}}_{i})}{p_{T-t}({\mathbf{x}}_{j})}\frac{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{i})}{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{j})}={\bm{Q}}_{t}^{[i,j]}\cdot\frac{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{i})}{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{j})}$ (49)
Similar ideas are applied to classifier guidance for discrete diffusion models [32], where the matrix ${\bm{R}}_{t}^{\mathbf{y}}=[\frac{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{i})}{p_{T-t}({\mathbf{y}}|{\mathbf{x}}_{j})}]_{i,j}$ is called a guidance rate matrix. We compute ${\bm{R}}_{t}^{\mathbf{y}}$ at ${\mathbf{x}}_{j}$-column by enumerating every neighboring ${\mathbf{x}}_{i}$ and calculating $\frac{p_{t}({\mathbf{y}}|{\mathbf{x}}_{i})}{p_{t}({\mathbf{y}}|{\mathbf{x}}_{j})}$ for each ${\mathbf{x}}_{i}$. The discrete version of DPS can be summarized by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}={\bm{Q}}_{t}{\bm{R}}_{t}^{\mathbf{y}}p_{T-t}.$ (50)
However, the discrete version of DPS is very time-consuming, especially when the vocabulary size is large, for it enumerates $(N-1)\times D$ number of neighboring ${\mathbf{x}}$ when computing ${\bm{R}}_{t}^{\mathbf{y}}$. We find it slow for binary MNIST ($N=2$) and DNA generation tasks $(N=4)$ but unaffordable for monophonic music generation $(N=129)$.
C.2.2 SVDD
SVDD [27] aims to sample from the distribution $p^{\beta}({\mathbf{x}}_{0})\ \ \propto\ \ p({\mathbf{x}}_{0})\exp(\beta r({\mathbf{x}}_{0}))$, which is equivalent to the regularized MDP problem:
$p^{\beta}({\mathbf{x}}_{0})=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{{\mathbf{x}}_{0}\sim\pi}r({\mathbf{x}}_{0})-\mathsf{KL}(\pi\|p)/\beta$ (51)
@@ C.2.2 SVDD
$v_{t}({\mathbf{x}}_{t})=\log\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})}[\exp(\beta r({\mathbf{x}}_{0}))]/\beta$ (52)
@@
$p_{t}^{\star}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})\ \ \propto\ \ p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})\exp(\beta v_{t}({\mathbf{x}}_{t})).$ (53)
In time step $t$, SVDD samples a batch of $M$ particles from the unconditional distribution $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})$, and conduct importance sampling according to $\exp(\beta v_{t}({\mathbf{x}}_{t}))$.
Although [27] is initially designed for guided diffusion generation; it also applies to solving inverse problems by carefully choosing reward functions. We consider $r({\mathbf{x}})=-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}}\|_{0}/\sigma_{\mathbf{y}}$ and $\beta=1$, so that it samples from the posterior distribution $p({\mathbf{x}}|{\mathbf{y}})$. As recommended in [27], we choose $\beta=\infty$ ($\alpha=0$ in their notation) in practice, so the importance sampling reduces to finding the particle with the maximal value in each iteration. We use SVDD-PM, a training-free method provided by [27] in our experiments. It approximates the value function by $v_{t}({\mathbf{x}}_{t})=r(\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t}))$, where $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ is an approximation of $\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{t}]$. In practice, we find that approximating $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ by Monte Carlo sampling with a few-step Euler sampler achieves slightly better results. We use $3$ Monte Carlo samples to estimate $v_{t}({\mathbf{x}}_{t})$ in our experiments.
C.2.3 SMC
Sequential Monte Carlo (SMC) methods evolve multiple particles to approximate a series of distributions, eventually converging to the target distribution. Specifically, in our experiments, we implement the SMC method to sample from $p({\mathbf{x}}_{t}|{\mathbf{y}})$ for $t=T,T-1\dots,0$, using the unconditional discrete diffusion sampler $p({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1})$ as the proposal function.
We maintain a batch of $M=20$ particles $\{{\mathbf{x}}^{[m]}\}$. At time $t$, we sample ${\mathbf{x}}_{t}^{[m]}\sim p({\mathbf{x}}_{t}|{\mathbf{x}}_{t+1}^{[m]})$ by the pretrained discrete diffusion model and estimate the likelihood $p({\mathbf{y}}|{\mathbf{x}}_{t}^{[m]})=\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0}|{\mathbf{x}}_{t}^{[m]})}p({\mathbf{y}}|{\mathbf{x}}_{0})$ by Monte Carlo sampling. We then resample the particles $\{{\mathbf{x}}_{t}^{[m]}\}$ according to their weights $w_{t}^{[m]}=p({\mathbf{y}}|{\mathbf{x}}_{t}^{[m]})/p({\mathbf{y}}|{\mathbf{x}}_{t+1}^{[m]})$. In practice, we find that we have to carefully tune a hyperparameter $\beta$, where $w_{t}^{\beta}$, or otherwise the resampling step can easily degenerate to $\arg\max$ or uniform random sampling.
C.3 Hyperparameters
We use an annealing noise schedule of $\eta_{k}=\eta_{\min}^{k/K}\eta_{\max}^{1-k/K}$ with $\eta_{\min}=10^{-4}$ and $\eta_{\max}=20$. We run SGDD for $K$ iterations. In each likelihood sampling step, we run Metropolis-Hastings for $T$ steps, while in each prior sampling step, we run a few-step Euler discrete diffusion sampler with $H$ steps. The hyperparameters used for each experiment are listed in Table 7. We also include the dimension of data spaces $\mathcal{X}^{D}$ for each experiment in Table 7, where $|\mathcal{X}|=N$.